a body is thrown verticallyuo from the ground passes the height 10.2 m twice at an interval of 10 s. what is initial velocity
 

Grade:11

1 Answers

gourav
4 years ago
now h=ut-1/2gt2    take g=10m/s2
h = 2 * 10.5=  21 m
u=?
t= 10s
21 = 10u-1/2*10*100
21=10u-500
521=10u
u=52.1m/s